OpenStudy (ria23):

Can someone check my answers for this question? A basketball is dropped from an original height of 10 feet. On each consecutive bounce, it's height reaches 3/4 the previous height. What is the total distance the ball will have traveled between the original drop and the 4th bounce? Show all your work. Don't forget the units.

OpenStudy (ria23):

\[S=\frac{ a1 }{(1-r)}\]\[S=\frac{ 7.5 }{ (1-.75) }\]\[S=30*2=60\]\[S=60+10\]\[S=70\]

OpenStudy (dmndlife24):

1st bounce: 3/4 of 10 = 7.5 ft. 2nd bounce: 3/4 of 7.5 = 5.625 ft. 3rd bounce: 3/4 of 5.625 = 4.219 ft. 4th bounce: 3/4 of 4.219 = 3.164 ft.

OpenStudy (ria23):

That's... A lot easier than what I did. >-< Thank yhu...

OpenStudy (dmndlife24):

Yw!